# Choosing observables that capture critical slowing down before tipping points: A Fokker-Planck operator approach

### v1

@@ III.1 Interpretation of Fokker-Planck eigenfunctions in one dimension
We first study the FP eigenfunctions in a one-dimensional double-well potential (DW1) with additive noise
@@
where $\beta$ is a control parameter. The potential is shown in Fig. 1e for different values of $\beta$ The deterministic drift of the system undergoes a saddle-node bifurcation at $\beta_{c}=\sqrt{8/27}\approx 0.544331$, where one of the potential wells disappears. From here on we will simply refer to this as bifurcation or TP, also when referring to the stochastic system. In Fig. 1 we show results where the forward eigenfunctions $\psi_{n}$ have been determined numerically, by a discrete approximation of $\mathcal{L}$ using the finite difference method by Chang and Cooper [39], and then an eigendecomposition of the obtained matrix using an implicitly restarted Arnoldi method (scipy.sparse.linalg.eigs package implementation of ARPACK).
@@
As $\beta$ is increased from zero towards the bifurcation, for low noise $\pi(\mathbf{x})$ quickly becomes heavily asymmetric with a dominant peak at the deeper potential well (Fig. 1a), and a peak around the shallower well that is orders of magnitude smaller. $\psi_{n}$ for small $n>0$ correspond to distinct patterns that modify the density such that it takes longest until statistical equilibrium is reached, given that the pattern projects significantly on the initial density $\rho_{0}$.
@@
Figure 1: Eigenfunctions of the Fokker-Planck operator for the one-dimensional double-well model (16) for a range of parameter values from $\beta=0.05$ to $\beta=0.53$ drawn with increasing contrast. The bifurcation occurs at $\beta\approx 0.544$. The eigenfunctions are estimated by eigendecomposition of the discrete approximation of the FP operator via the scheme of Chang and Cooper [39]. a-d The first four eigenfunctions of the forward FP operator for $\sigma=0.25$. e Associated potential $V(x)$. f-h The first three non-trivial eigenfunctions of the backward operator $\phi_{1,2,3}$ (rescaled for each parameter value to have a maximum value of 1). The green dots in g indicate the locations of the saddle point for the respective parameter values. The vertical red line is the inflection point of the potential in the shallower well, which is independent of $\beta$.
@@
The slowest decaying pattern $\psi_{1}$ describes the situation where the probability mass in one of the wells is initially larger as it should be according to $\pi(\mathbf{x})$ (Fig. 1b). For the equilibration of such a configuration, part of the probability mass needs to diffuse uphill and overcome the potential barrier. In metastable systems with low noise, as considered here, this is associated with a long time scale, and $\lambda_{1}$ is approximately equal to the escape rate out of the shallow well.
@@
The pattern $\psi_{2}$ gives a large contribution when the initial density is concentrated more prominently in the vicinity of the saddle (as compared to $\pi(\mathbf{x})$). The function shows minima slightly outwards (larger $|x|$)) of the two stable fixed points, and a broad double maximum around the saddle point (Fig. 1c). This pattern can be interpreted as additional mass that survives outside the vicinity of the two minima for some time $\lambda_{2}^{-1}$ due to the asymmetry of each well, i.e., the smaller curvature of the potential towards the saddle. In other words, the relaxation towards equilibrium is slower in the vicinity of the saddle and on the sides of the wells that are facing the saddle. As the bifurcation is approached, the segment of the potential within the shallow well that faces the saddle becomes smaller, and thus this is the relevant mode that carries the CSD. In the one-dimensional case, higher eigenfunctions are less important for our analysis, representing higher-order corrections (see Fig. 1d for $\psi_{3}$).
@@
While $\phi_{0}$ can be called the trivial eigenfunction because it is constant, and $\phi_{1}$ the dominant eigenfunction since $|\lambda_{2}-\lambda_{1}|\gg|\lambda_{1}|$, we refer to $\phi_{2}$ as the first subdominant eigenfunction. $\phi_{2}$ peaks close to the saddle point, and converges to 0 at the deep well while reaching lower values in the shallow well. On time scales of order $\lambda_{2}^{-1}$, expectation values are thus altered when starting close to the saddle. $\phi_{3}$ (and similarly higher eigenfunctions) is noteworthy in the sense that it is not a monotonic function within the shallow well. It first increases towards the inflection point in the shallow well (Fig. 1h), and then decreases again towards the saddle. In contrast, for the purpose of EWS, we are interested in observables that are monotonic from the base attractor towards the edge state, because otherwise fluctuations of increasing length along the critical d.o.f towards the edge state due to CSD are suppressed in the measured observable. ¨
The CSD as $\beta$ is changed towards the bifurcation is reflected in the eigenvalues, though this requires a low noise level to be observed. Since $\psi_{2}$ captures the slowing of relaxation towards equilibrium as the curvature in the shallow well decreases, the relaxation rate $\lambda_{2}$ should go towards zero as a manifestation of CSD. This can be seen for low noise levels in Fig. 2b,c. For finite noise levels the relaxation time remains finite (Fig. 2a), and even slightly decreases very close to the bifurcation. It is bounded by the noise-induced escape rate $\lambda_{1}$, which decreases drastically and becomes O(1) at the bifurcation, at which point the potential is so flat that the relevant time scale for the decay of $\psi_{2}$ is not the deterministic relaxation, but pure diffusion dynamics. Additionally, the decrease in distance of the saddle and the fixed point in the shallow well may play a role.
Figure 2: a-c Eigenvalues $\{\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}\}$ of the FP operator for the one-dimensional double-well potential as function of the control parameter $\beta$, for different noise levels $\sigma$. The critical value corresponding to the bifurcation is marked by the vertical dashed line. d Eigenvalues $\{\lambda_{1},\lambda_{2},...,\lambda_{7}\}$ of the FP operator of the two-dimensional DW2 model (19) as a function of the control parameter $e$, using the noise level $\sigma=0.3$. The bifurcation point is marked with the vertical dashed line.
@@ III.2 Eigenfunctions in two dimensions
In the previous one-dimensional example, there were no other slow d.o.f that compete with the critical d.o.f to be in the position of mode $\phi_{2}$. But generally the situation is different, especially in high-dimensional systems with multi-scale behaviour, where the correct physical mode first needs to slow down enough so that it emerges as $\phi_{2}$. We illustrate this with a system of two variables $x$ and $y$ in the double-well potential (DW2)
@@
3 Adding Gaussian white noise independently to both variables yields the system of stochastic differential equations
@@
where $W_{x,t}$ and $W_{y,t}$ are independent, standard Wiener processes. Fixed values $a=2.5$, $b=0.5$, $c=0.2$, $d=0.5$ are used, and $e$ is the control parameter. For small $e$, there are two stable fixed points and one saddle point in the deterministic system. There is a saddle-node bifurcation of the deterministic drift at $e_{c}\approx 1.73$, where the potential well with $x>0$ disappears. Figure 3f shows isolines of the potential, as well as the fixed points and basin boundary at $e=0.5$.
The first non-trivial eigenfunction $\psi_{1}$ is again related to the slow transport of density from one well to the other (Fig. 3b), and accordingly the backward function $\phi_{1}$ is approximately constant in the two basins, with a steep transition layer along the basin boundary (Fig. 3g). Next, compared to DW1 there is an additional eigenmode because of the slow time scale from the generally slower deterministic dynamics in the $y$ direction. Far from the bifurcation, $\psi_{2}$ represents probability mass that is more slowly contracted in the $y$-direction and for some time ($\lambda_{2}^{-1}$) has the tendency to linger at strongly negative $y$ values (Fig. 3c) instead of converging to either fixed point. Hence, $\phi_{2}$ identifies initial conditions that take longest to converge to either of the two wells along the $y$-direction (Fig. 3h). The next mode $\psi_{3}$ corresponds (at this value of the control parameter $e$) to $\psi_{2}$ of the one-dimensional case, i.e., a result of slow convergence to the fixed points in the more flat parts of the asymmetric potential wells towards the saddle point (Fig. 3d). Correspondingly, $\phi_{3}$ peaks near the saddle, and it shows that in particular initial conditions starting near the stable manifold of the saddle will lead to a slow relaxation of conditional expectation values. $\psi_{4}$ is a higher-order, antisymmetric pattern, analogous to $\psi_{3}$ of the DW1 model.
@@
As $e$ approaches the bifurcation, and given the noise is low enough, the eigenvalues of the abovementioned patterns cross (Fig. 2d), leading to a different ordering of the modes (Fig. 3k-t). The eigenvalue of the pattern associated with the d.o.f. in the $y$-direction decreases (gray band in Fig. 2d), and the pattern drops to higher $n$. Instead, the pattern related to the low potential curvature towards the saddle (yellow band in Fig. 2d) becomes $\phi_{2}$ (Fig. 3r), i.e., the pattern subdominant only to the pattern $\phi_{1}$ that reflects noise-induced escape.
@@ III.3 Eigenfunction reconstruction from diffusion maps and observables for early-warning signals
We now reconstruct the backward eigenfunctions $\phi_{n}(x)$ via the DM approach from data of the DW2 model (19) obtained by simulation with an Euler-Maruyama scheme with time step $dt=0.005$. We simulate an ensemble of 100 uniformly distributed initial conditions covering both wells for a fixed simulation time $T=100$, allowing the ensemble to converge to $\pi(\mathbf{x})$. We only use data after $t=75$, i.e., we discard any transient dynamics. Finally, the simulated data is subsampled and all ensemble members combined to yield a set of 15,000 data points.
Figure 4: a-d First four diffusion coordinates $\xi_{1,2,3,4}$, which approximate the eigenfunctions $\phi_{1,2,3,4}$ of $\mathcal{L}^{\star}$, of the DW2 model (19) with control parameter $e=0.5$ and noise strength $\sigma=0.6$, estimated by DM on simulated data sampling the whole phase space. The diffusion coordinates are evaluated at evenly spaced grid points using Eq. 14-15. e,f First two non-trivial diffusion coordinates of the DW2 model (19) from simulated data restricted to dynamics that remains in the shallow well, with control parameter $e=0.5$ and noise strength $\sigma=0.3$.
The scaled eigenvectors $\xi_{n}$ (i.e. the diffusion coordinates) obtained from a spectral decomposition of the Markov matrix (13) define functions that can be evaluated approximately at any point in the original phase space via Eq. 14-15. Evaluation on an evenly spaced grid for DW2 shows that the first few non-trivial diffusion coordinates $\xi_{1,2,3,4}$ are indeed in good qualitative agreement with the (scaled) eigenfunctions $\phi_{1,2,3,4}$ obtained from the discretized FP operator (compare Fig. 4a-d and Fig. 3g-j).
We now restrict our attention to the scenario of bifurcation-induced tipping, where we assume the system resides in one of the metastable sets, and where we consider time scales much shorter than $\lambda_{1}^{-1}$ associated with noise-induced escape. We consider the well that contains the base state with $x>0$ (cf. Fig. 4 or Fig. 3). Assuming a slowly varying control parameter, the system is close to a quasi-stationary distribution $p_{qs}(x)\approx\psi_{0}(x)+c_{1}\psi_{1}(x)$ at any given instantaneous control parameter value. Here, $c_{1}\psi_{1}$ compensates $\psi_{0}$ such that all mass is concentrated in the shallow well. The dominant eigenfunction $\phi_{1}$ is approximately constant in the shallow well, and thus variations in expectation values are determined by $\phi_{2}$ and onward. When sufficiently close to the bifurcation, the first backward eigenfunctions are almost constant $\phi_{2}\approx\phi_{3}\approx\phi_{4}\approx 0$ in the basin of the alternative state (Fig. 3q-t). Hence, it should be possible to approximate them from data restricted to the basin of the base state.
@@
Figure 5: a, d Simulated data points of the DW2 model (19) in the space spanned by the first three diffusion coordinates, for parameter value $e=0.5$ far from the bifurcation (left) and for $e=1.5$ (right), which is closer to the bifurcation at $e\approx 1.73$. The lower panels b,c,e,f show the same data in two-dimensional projections onto the diffusion coordinates ($\xi_{1}$,$\xi_{2}$) and ($\xi_{1}$,$\xi_{3}$). We use a lower noise level of $\sigma=0.09$ in order to obtain simulation data restricted to the shallow well when very close to the bifurcation, and thus the relation of $\xi_{1}$ and $\xi_{2}$ is different compared to Fig. 4e-f, where $\sigma=0.3$.
@@
The leading $\xi_{n}$ are not necessarily all independent d.o.f’s. In the system restricted to the shallow well, the critical d.o.f becomes the slowest upon approaching the bifurcation and the time scale separation with respect to all other d.o.f’s becomes larger, at which point the first few backward eigenfunctions (and hence the first few diffusion coordinates) are all expected to parameterize the slowest d.o.f. This is particular for single-well systems. For instance, in a multi-dimensional parabolic potential with a slow variable $x$ and a spectral gap, $\xi_{1}$ is a function of $x$ and the next $k$ eigenfunctions $\xi_{k}$ (with $k$ dependent on the magnitude of the spectral gap) are polynomially related to $\xi_{1}$ [32, 36]. In this case, $\xi_{1}$ is sufficient as a reduction coordinate and is the only diffusion coordinate that indicates monotonically how far fluctuations evolve towards the saddle.
Accordingly, in the DW2 example restricted to the shallow well, the dynamics in the space of the diffusion coordinates $\xi_{1,2,3}$ evolves from a two-dimensional to a one-dimensional manifold as the bifurcation is approached (Fig. 5).
Further away from the bifurcation, $\xi_{2}$ represents the slow mode in the $y$-direction, and the two other diffusion coordinates are quadratically related and parameterize the d.o.f related to the asymmetry of the potential well towards the edge state (Fig. 5a-c). Close to the bifurcation, the dynamics becomes constrained to an approximately one-dimensional curve where $\xi_{2}$ ($\xi_{3}$) is a quadratic (cubic) function of $\xi_{1}$ (Fig. 5d-f). The observed pattern formerly associated with $\xi_{2}$ at $e=0.5$ drops to higher eigenfunctions.
@@
Figure 6: a Distribution of the $\xi_{1}$-values evaluated at the data points of a DW2 simulation (19) at $e=1.5$ and $\sigma=0.09$ (red line). The black dashed line is a distribution of the same $\xi_{1}$ (i.e. obtained from the DM at $e=1.5$), but evaluated on data points of a simulation at $e=0.5$ via the Nyström interpolation (14)-(15). b-d The diffusion coordinate $\xi_{1}$, estimated for DW2 simulation data close to the tipping point ($e=1.65$), is fit to a polynomial of the state variables (x,y), and used as observable to detect CSD by evaluating it on residual data of simulations at lower values of the control parameter. All simulations are initialized in the shallow well. Shown are distributions of the values of a linear (d) and cubic (c) polynomial, normalized to the fluctuations at $e=0.2$. Panel b shows the variance increase of the linear and non-linear observable, as well as the observable $x-y$, compared to the variance of the fluctuations at $e=0.25$. e-g Polynomial fits of the first non-trivial diffusion coordinate $\xi_{1}$ estimated from simulations of the DW2 system (19) (dots) with $\sigma=0.09$. Panel e is a cubic fit to data at the parameter value $e=0.5$ far from the bifurcation, and (f,g) is a cubic and linear fit for $e=1.5$, which is closer to the bifurcation.
@@
The observations above suggest three main ways to leverage information contained in the diffusion coordinates $\xi_{n}$ for early-warning of TPs. First, one may observe the qualitative change of the functional dependencies of the first $\xi_{n}$ as a result of the emerging time scale separation, as just discussed. Second, one can directly evaluate and compare $\xi_{1}$ for data sets obtained at different observational time slices (i.e. for different values of the control parameter) via the Nyström interpolation (14)-(15). In particular, we can estimate $\xi_{1}$ from a data set believed to be closest to a TP, for example from climate observations closest to present-day, and then evaluate the observable $\xi_{1}$ on data sets sampled further away from the TP, for example using climate observations of the past. If the variability and correlation of the values of $\xi_{1}$ is increased significantly in the former data set, this is an indication of decrease in resilience in the critical d.o.f and an impending TP. Note that the same normalization of the variables that is applied before the DM algorithm to the data set where $\xi_{1}$ is estimated has to be applied to data at parameter values further away from the bifurcation. How $\xi_{1}$ can be used in this manner is illustrated in Fig. 6a, where a clear change in variability is seen when evaluating $\xi_{1}$ estimated at $e=1.5$ on data simulated farther away from the bifurcation at $e=0.5$. Note that here and in most of the following we focus our presentation on the variance as EWS, but similar plots could be shown for the autocorrelation.

### v2

@@ III.1 Interpretation of Fokker-Planck eigenfunctions in one dimension
We first study the FP eigenfunctions in a one-dimensional system with additive noise
@@
where $\beta$ is a control parameter. The potential is shown in Fig. 1e for different values of $\beta$ . The deterministic drift of the system undergoes a saddle-node bifurcation at $\beta_{c}=\sqrt{8/27}\approx 0.544331$, where one of the potential wells disappears. From here on we will simply refer to this as bifurcation or TP, also when referring to the stochastic system. We determine the forward eigenfunctions $\psi_{n}$ numerically by a discrete approximation of $\mathcal{L}$ using the finite difference method by Chang and Cooper [47], and then an eigendecomposition of the obtained matrix using an implicitly restarted Arnoldi method (scipy.sparse.linalg.eigs package implementation of ARPACK). Since this is a gradient system the forward and backward eigenfunctions are related by $\phi_{n}(\mathbf{x})=\pi(\mathbf{x})^{-1}\psi_{n}(\mathbf{x})$, with the same, real $\lambda_{n}$. Hence, we can determine both $\psi_{n}$ and $\phi_{n}$, which is shown in Fig. 1.
@@
As $\beta$ is increased from zero towards the bifurcation, for low noise $\pi(\mathbf{x})$ quickly becomes heavily asymmetric with a dominant peak at the deeper potential well (Fig. 1a), and a peak around the shallower well that is orders of magnitude smaller. The $\psi_{n}$ for small $n>0$ correspond to distinct patterns that modify the density such that it takes longest until statistical equilibrium is reached, given that the initial density $\pi_{0}(\mathbf{x})$ projects significantly on those pattern.
@@
Figure 1: Eigenfunctions of the Fokker-Planck operator for the one-dimensional double-well model (16) for a range of parameter values from $\beta=0.05$ to $\beta=0.53$ drawn with increasing contrast. The bifurcation occurs at $\beta\approx 0.544$. The eigenfunctions are estimated by eigendecomposition of the discrete approximation of the FP operator via the scheme of Chang and Cooper [47]. a-d The first four eigenfunctions of the forward FP operator for $\sigma=0.25$. e Associated potential $V(x)$. f-h The first three non-trivial eigenfunctions of the backward operator $\phi_{1,2,3}$ (rescaled for each parameter value to have a maximum value of 1). The green dots in g indicate the locations of the saddle point for the respective parameter values. The vertical dashed line is the inflection point of the potential in the shallower well, which is independent of $\beta$.
@@
The slowest decaying pattern $\psi_{1}$ describes the situation where the probability mass in one of the wells is initially larger as it should be according to $\pi(\mathbf{x})$ (Fig. 1b). For the equilibration of such a configuration, part of the probability mass needs to diffuse uphill and overcome the potential barrier. In systems with low noise this is associated with a long time scale, and $|\lambda_{1}|$ is approximately equal to the escape rate out of the shallow well.
@@
The pattern $\psi_{2}$ gives a large contribution when the initial density is concentrated more prominently in the vicinity of the saddle (as compared to $\pi(\mathbf{x})$). The function shows minima slightly outwards (larger $|x|$)) of the two stable fixed points, and a broad double maximum around the saddle point (Fig. 1c). This pattern can be interpreted as additional mass that survives outside the vicinity of the two minima for some time $|\lambda_{2}|^{-1}$ due to the asymmetry of each well, i.e., the smaller curvature of the potential towards the saddle. In other words, the relaxation towards equilibrium is slower in the vicinity of the saddle and on the sides of the wells that are facing the saddle. As the bifurcation is approached, the segment of the potential within the shallow well that faces the saddle becomes more flat, and thus $\psi_{2}$ is the relevant mode carrying CSD. In the one-dimensional case, higher eigenfunctions are less important for our analysis, representing higher-order corrections (see Fig. 1d for $\psi_{3}$).
@@
While the constant $\phi_{0}$ can be called the “trivial” eigenfunction, and $\phi_{1}$ the dominant eigenfunction since $|\lambda_{2}-\lambda_{1}|\gg|\lambda_{1}|$, we refer to $\phi_{2}$ as the first subdominant eigenfunction. $\phi_{2}$ peaks close to the saddle point, and converges to 0 at the deep well while reaching lower values in the shallow well. On time scales of order $|\lambda_{2}|^{-1}$, expectation values are thus altered when starting close to the saddle. $\phi_{3}$ (and similarly higher eigenfunctions) is non-monotonic within the shallow well. It first increases towards the inflection point in the shallow well (Fig. 1h), and then decreases again towards the saddle. For the purpose of EWS, we are, however, interested in observables that are monotonic from the base attractor towards the edge state, because otherwise fluctuations of increasing length along the critical d.o.f towards the edge state due to CSD are suppressed in the measured observable.
Depending on the noise level, the CSD as $\beta$ is changed towards the bifurcation may be reflected in the eigenvalues. Since $\psi_{2}$ captures the slowing of relaxation towards equilibrium as the curvature in the shallow well decreases, the relaxation rate $|\lambda_{2}|$ should go towards zero as a manifestation of CSD. This can indeed be seen for low noise levels in Fig. 2b,c. But for finite noise levels $|\lambda_{2}|$ remains finite (Fig. 2a), and even slightly increases close to the bifurcation. It is bounded by the noise-induced escape rate $|\lambda_{1}|$, which increases drastically and becomes O(1) at the bifurcation, at which point the potential is so flat that the relevant time scale for the decay of $\psi_{2}$ is not set by deterministic relaxation, but pure diffusion dynamics. Additionally, the approach of saddle and fixed point in the shallow well may play a role.
Figure 2: a-c Eigenvalues $\{\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}\}$ of the FP operator for the one-dimensional double-well potential as function of the control parameter $\beta$, for different noise levels $\sigma$. The critical value corresponding to the bifurcation is marked by the vertical dashed line. d Eigenvalues $\{\lambda_{1},\lambda_{2},...,\lambda_{7}\}$ of the FP operator of the two-dimensional double-well (19) as a function of the control parameter $e$, using the noise level $\sigma=0.3$. The bifurcation point is marked with the vertical dashed line.
@@ III.2 Eigenfunctions in two dimensions
In the previous one-dimensional example, there were no other slow d.o.f. that compete with the critical d.o.f. to be in the position of mode $\phi_{2}$. But generally the situation is different, especially in high-dimensional systems with multi-scale behaviour, where the correct physical mode first needs to slow down enough so that it emerges as $\phi_{2}$. We illustrate this with a system of two variables $x$ and $y$ in the double-well potential
@@
Adding Gaussian white noise independently to both variables yields the system
@@
where $W_{x,t}$ and $W_{y,t}$ are independent, standard Wiener processes. $a=2.5$, $b=0.5$, $c=0.2$, $d=0.5$ are chosen, and $e$ is the control parameter. For small $e$, there are two stable fixed points and one saddle point in the deterministic system. There is a saddle-node bifurcation of the deterministic drift at $e_{c}\approx 1.73$, where the potential well with $x>0$ disappears. Figure 3f shows isolines of the potential, as well as the fixed points and basin boundary at $e=0.5$.
The first non-trivial eigenfunction $\psi_{1}$ is again related to the slow transport of density from one well to the other (Fig. 3b), and accordingly the backward function $\phi_{1}$ is approximately constant in the two basins (Fig. 3g). Next, compared to the 1-d double well there is an additional eigenmode because of the slow time scale from the generally slower deterministic dynamics in the $y$ direction. Far from the bifurcation, $\psi_{2}$ represents probability mass that is more slowly contracted in the $y$-direction and for some time ($|\lambda_{2}|^{-1}$) has the tendency to linger at strongly negative $y$ values (Fig. 3c) instead of converging to either fixed point. Hence, $\phi_{2}$ identifies initial conditions that take longest to converge to either of the two wells along the $y$-direction (Fig. 3h). The next mode $\psi_{3}$ corresponds (at this value of the control parameter $e$) to $\psi_{2}$ of the one-dimensional case, i.e., a result of slow convergence to the fixed points in the more flat parts of the asymmetric potential wells towards the saddle point (Fig. 3d). Correspondingly, $\phi_{3}$ peaks near the saddle, and it shows that in particular initial conditions starting near the stable manifold of the saddle will lead to a slow relaxation of conditional expectation values. $\psi_{4}$ is a higher-order, antisymmetric pattern, analogous to $\psi_{3}$ of the 1-d double well.
@@
As $e$ approaches the bifurcation, the eigenvalues of the abovementioned patterns cross (Fig. 2d), leading to a different ordering of the modes (Fig. 3k-t). The eigenvalue of the pattern associated with the d.o.f. in the $y$-direction decreases (gray band in Fig. 2d), and the pattern drops to higher $n$. Instead, the pattern related to the low potential curvature towards the saddle (yellow band in Fig. 2d) becomes $\phi_{2}$ (Fig. 3r), i.e., the pattern subdominant only to the pattern $\phi_{1}$ that reflects noise-induced escape. In order for the critical d.o.f. to reach the position of $\phi_{2}$ before a transition, it is crucial that the noise is low enough. If this is not the case and there is a very slow (and unchanging) d.o.f, the moment when the critical d.o.f gets slow enough, i.e., when $|\lambda_{3}|$ approaches $|\lambda_{2}|$, may only occur at a parameter value where noise-induced escape is already common, i.e., when $|\lambda_{1}|$ approaches $|\lambda_{2}|$ and $|\lambda_{3}|$ from below and reaches the same order of magnitude.
@@ III.3 Eigenfunction reconstruction from diffusion maps and observables for early-warning signals
We now reconstruct $\phi_{n}(x)$ via the DM approach from data of the 2-d double-well (19) obtained by simulation with an Euler-Maruyama scheme with time step $dt=0.005$. We simulate an ensemble of 100 uniformly distributed initial conditions covering both wells for a fixed simulation time $T=100$, allowing the ensemble to converge to $\pi(\mathbf{x})$. We only use data after $t=75$, i.e., we discard any transient dynamics. Finally, the simulated data is subsampled and all ensemble members combined to yield a set of 15,000 data points.
Figure 4: Reconstruction of eigenfunctions of the 2-d double-well (19) using the DM algorithm. a-d First four diffusion coordinates $\xi_{1,2,3,4}$ that approximate $\phi_{1,2,3,4}$, obtained at control parameter $e=0.5$ and noise strength $\sigma=0.6$, and estimated by DM on simulated data sampling the whole phase space. The diffusion coordinates are evaluated at evenly spaced grid points using Eq. 14-15. e,f First two non-trivial diffusion coordinates of (19) from simulated data restricted to dynamics that remains in the shallow well, with control parameter $e=0.5$ and noise strength $\sigma=0.3$.
The scaled eigenvectors $\xi_{n}$ (i.e. the diffusion coordinates) obtained from a spectral decomposition of the Markov matrix (13) define functions that can be evaluated approximately at any point in the original phase space via Eq. 14-15. Evaluation on an evenly spaced grid for 2-d double-well shows that the first few non-trivial diffusion coordinates $\xi_{1,2,3,4}$ are indeed in good qualitative agreement with the (scaled) eigenfunctions $\phi_{1,2,3,4}$ obtained from the discretized FP operator (compare Fig. 4a-d and Fig. 3g-j).
We now restrict our attention to the scenario of bifurcation-induced tipping, where the system resides in one of the metastable sets, and where we consider time scales much shorter than $|\lambda_{1}|^{-1}$ associated with noise-induced escape. We consider the well that contains the base state with $x>0$ (cf. Fig. 4 or Fig. 3). Assuming a slowly varying control parameter, the system relaxes towards a quasi-stationary distribution $p_{qs}(x)\approx\psi_{0}(x)+c_{1}\psi_{1}(x)$ at any given instantaneous control parameter value. Here, $c_{1}\psi_{1}$ compensates $\psi_{0}$ such that all mass is concentrated in the shallow well. The dominant eigenfunction $\phi_{1}$ is approximately constant in the shallow well, and thus variations in expectation values are determined by $\phi_{2}$ and onward. When sufficiently close to the bifurcation, the first backward eigenfunctions are almost constant $\phi_{2}\approx\phi_{3}\approx\phi_{4}\approx 0$ in the basin of the alternative state (Fig. 3q-t). Hence, it should be possible to approximate them from data restricted to the basin of the base state.
@@
Figure 5: a, d Simulated data points of the 2-d double-well (19) in the space spanned by the first three diffusion coordinates, for parameter value $e=0.5$ far from the bifurcation (left) and for $e=1.5$ (right), which is closer to the bifurcation at $e\approx 1.73$. The lower panels b,c,e,f show the same data in two-dimensional projections onto the diffusion coordinates ($\xi_{1}$$\xi_{2}$) and ($\xi_{1}$$\xi_{3}$). We use a lower noise level of $\sigma=0.09$ in order to obtain simulation data restricted to the shallow well when very close to the bifurcation, and thus the relation of $\xi_{1}$ and $\xi_{2}$ is different compared to Fig. 4e-f, where $\sigma=0.3$.
@@
The leading $\xi_{n}$ are not necessarily all independent d.o.f’s. In the system restricted to the shallow well, the critical d.o.f becomes the slowest upon approaching the bifurcation and the time scale separation with respect to all other d.o.f’s becomes larger, at which point the first few backward eigenfunctions (and hence the first few diffusion coordinates) are all expected to parameterize the slowest d.o.f. This is particular for single-well systems. For instance, in a multi-dimensional parabolic potential with a slow variable $x$ and a spectral gap, $\xi_{1}$ is a function of $x$ and the next $k$ eigenfunctions $\xi_{k}$ (with $k$ dependent on the magnitude of the spectral gap) are polynomially related to $\xi_{1}$ [32, 42]. In this case, $\xi_{1}$ is sufficient as a reduction coordinate and is the only diffusion coordinate that indicates monotonically how far fluctuations evolve towards the saddle. Accordingly, in the 2-d double-well restricted to the shallow well, the dynamics in the space of the diffusion coordinates $\xi_{1,2,3}$ evolves from a two-dimensional to a one-dimensional manifold as the bifurcation is approached (Fig. 5).
@@
Figure 6: Observables for EWS in the 2-d double well model (19). a Distribution of the $\xi_{1}$-values evaluated at simulated data points using $e=1.5$ and $\sigma=0.09$ (red line). The black dashed line is a distribution of the same $\xi_{1}$ (i.e. obtained from the DM at $e=1.5$), but evaluated on data points simulated at $e=0.5$ via the Nyström interpolation (14)-(15). b-d The diffusion coordinate $\xi_{1}$, estimated for simulation data close to the tipping point ($e=1.65$), is fit to a polynomial of the state variables (x,y), and used as observable to detect CSD by evaluating it on residual data of simulations at lower values of the control parameter e. All simulations are initialized in the shallow well. Shown are distributions of the values of a linear (d) and cubic (c) polynomial, normalized to the fluctuations at $e=0.2$. Panel b shows the variance increase of the linear and non-linear observable, as well as the observable xyx-y, normalized to the variance of the fluctuations at $e=0.25$. e-g Polynomial fits of $\xi_{1}$ estimated from simulations (dots) with $\sigma=0.09$. Panel e is a cubic fit to data at the parameter value $e=0.5$ far from the bifurcation, and (f,g) is a cubic and linear fit for $e=1.5$, which is closer to the bifurcation.
@@
The observations above suggest three main ways to leverage information contained in the diffusion coordinates $\xi_{n}$ for early-warning of TPs. First, one may observe the qualitative change of the functional dependencies of the first $\xi_{n}$ as a result of the emerging time scale separation, as just discussed. Second, one can directly evaluate and compare $\xi_{1}$ for data sets obtained at different observational time slices (i.e. for different values of the control parameter) via the Nyström interpolation (14)-(15). In particular, we can estimate $\xi_{1}$ from a data set believed to be closest to a TP, for example from climate observations closest to present-day, and then evaluate the observable $\xi_{1}$ on data sets sampled further away from the TP, for example using climate observations of the past. If the variability and correlation of the values of $\xi_{1}$ is increased significantly in the former data set, this is an indication of decrease in resilience in the critical d.o.f and an impending TP. This is illustrated in Fig. 6a, where a clear change in variability is seen when evaluating $\xi_{1}$ estimated at $e=1.5$ on data simulated farther away from the bifurcation at $e=0.5$. Note that the same normalization of the variables that is applied before the DM algorithm to the data set where $\xi_{1}$ is estimated has to be applied to data at parameter values further away from the bifurcation. Here and in most of the following we focus our presentation on the variance as EWS, but similar plots could be shown for the autocorrelation.